# Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization

### v2

@@ 4.1.2 Models Used
Large Language Models (LLMs):
For the code inference stage, we experiment with GPT-4o mini Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) to generate candidate visualization code from the reasoning paths. Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with previous work Yang et al. (2024). To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) as our visualization feedback model, which provides high-quality reference assessments.
Vision-Language Models (VLMs):
In order to assess the visual quality and correctness of the rendered plots, we incorporate vision evaluation models into our framework. Specifically, GPT-4o Achiam et al. (2023) is employed for detailed plot evaluation in all evaluation tasks. This setup ensures the thorough evaluation of both the syntactic correctness of the code and the aesthetic quality of the resulting visualizations.
@@ 4.1.3 Evaluation Metrics
In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability. MatPlotBench Bisong and Bisong (2019) assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions. Qwen-Agent Code Interpreter benchmark [^2] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md evaluates visualization models based on Visualization-Hard and Visualization-Easy. Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric. Previous studies showed that GPT-based VLM evaluations align well with human assessments Yang et al. (2024), hence VLM was used for evaluation.
@@
CoT Prompting Wei et al. (2022) 61.95 50 57.50 40.00 48.75
Chat2VIS Maddigan and Susnjak (2023) 56.98 53 59.36 36.50 47.93
MatPlotAgent Yang et al. (2024) 63.90 58 67.50 53.25 60.38
@@
CoT Prompting Wei et al. (2022) 53.56 61 40.00 63.89 51.95
Chat2VIS Maddigan and Susnjak (2023) 54.89 55 59.36 56.50 57.93
MatPlotAgent Yang et al. (2024) 56.31 58 77.62 51.50 64.56
@@
Table 1: Performance comparison of various methods across different benchmarks. Zero-Shot refers to directly generating code. CoT Prompting utilizes Chain of Thought Prompting. Visualization-Hard and Visualization-Easy refer to the Accuracy of Code Execution Results on different subsets of the Qwen-Agent Code Interpreter benchmark. Bold text indicates the best performance, underlined text indicates the second-best performance. denotes our proposed method.
4.1.4 Baseline Methods
We compare VisPath against competitive baselines: (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses Chain-of-Thought (CoT) prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al. (2024) first expands the query and then refines the code via a self-debugging loop with feedback. Moreover, our proposed framework VisPath generates three reasoning paths with corresponding visual feedback to refine the final output. [^3] Prompts are detailed in Appendix A. For a fair comparison aligned with our experimental setting, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well.
@@ 4.2 Experimental Analysis
VisPath is evaluated against four baselines: Zero Shot prompting, CoT prompting, Chat2VIS, and MatPlotAgent. The evaluation is conducted on MatPlotBench and the Qwen-Agent Code Interpreter benchmark using GPT-4o mini and Gemini 2.0 Flash, as shown in Table 1.
Zero-Shot prompting generates visualization code directly from natural language queries without intermediate reasoning. While computationally efficient, it often struggles to handle ambiguity or under-specification, resulting in misaligned or incomplete outputs. On MatPlotBench (GPT-4o mini), it achieves a Plot Score of 62.38 and an Executable Rate of 53%. CoT prompting further introduces a single reasoning step to expose intermediate decisions and improve interpretability. However, on MatPlotBench, it slightly underperforms Zero-Shot in both Plot Score and Executable Rate, indicating reliance on a fixed reasoning path may reduce adaptability to diverse input structures.
Chat2VIS extends CoT prompting by adopting prefix templates to improve coherence and reduce ambiguity in user instructions. While this approach is effective for well-structured or common query formats, its dependence on fixed templates limits adaptability when processing loosely specified or novel queries. Such limitation is evident in its performance on MatPlotBench, where it achieves a Plot Score of 56.98 and an Executable Rate of 53%. Furthermore, MatPlotAgent incorporates query expansion and iterative self-debugging mechanisms to enhance robustness. While effective at correcting execution-level errors, its revisions are confined to localized adjustments and do not address higher-order semantic ambiguities.
@@
Evaluated across multiple benchmark settings, VisPath notably outperforms baselines, achieving up to 9.14 point gains in Plot Score and a 10% point increase in Executable Rate. These improvements well demonstrate VisPath’s robustness in exploring diverse reasoning paths and iteratively refine outputs through structured visual feedback, effectively reducing semantic ambiguity and improving execution reliability.
@@ 4.3 Ablation Study
To further examine the robustness and design choices of VisPath, we conduct a series of ablation experiments in this section. Specifically, we analyze the following three aspects: (i) varying the number of generated reasoning paths, (ii) the effect of removing visual feedback during integration, and (iii) the contribution of visual feedback beyond binary executability.
@@ 4.3.1 Varying the Number of Reasoning Paths
To investigate the contribution of reasoning path diversity, we conducted ablation experiments by varying the number of generated reasoning paths $K$. In particular, we extended the range of $K$ from 2 to 8 to examine the effect of increased path multiplicity on the overall performance of VisPath, as shown in Figure 3.
We observe a consistent pattern across all model and dataset combinations: performance improves as $K$ increases from 2 to 3, confirming that limited diversity $(K=2)$ often fails to capture nuanced interpretations of user queries. For example, with GPT-4o mini on MatPlotBench, the Plot Score improves from 64.02 to 66.12 (+2.10), and the Executable Rate improves from 58% to 60% (+2 points). On Gemini 2.0 Flash, the Plot Score increases from 56.59 to 59.37 (+2.78), and the Executable Rate from 56% to 61% (+5 points).
While $K=4$ achieves the highest executable rate on MatPlotBench with GPT-4o mini (62%), we further extend our analysis up to $K$ = 8 to comprehensively assess the impact of reasoning path diversity. However, beyond $K=4$, we observe diminishing returns and even performance degradation, which is likely due to noisy or redundant reasoning paths. While added diversity initially aids interpretation, excessive expansion burdens the integration process and reduces overall efficiency.
@@ 4.3.2 Robustness with a Simple Integration
Model Feedback MatPlotBench Qwen-Agent
Plot Score Executable Rate (%) Avg.
GPT-4o mini (w/o) feedback 63.76 56 58.00
(w) feedback 66.12 60 63.96
Gemini 2.0 Flash (w/o) feedback 55.28 57 64.03
(w) feedback 59.37 63 68.98
Table 2: Performance comparison of VisPath with and without visual feedback. The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two LLMs.
Instead of refining each candidate visualization with feedback from vision-language feedback, this approach aggregates three candidate codes: each derived from a distinct reasoning path, without intermediate corrections. This setup reduces computational overhead and execution time while preserving the benefits of interpretive diversity.
As shown in Table 2, even under this simplified configuration, VisPath outperforms all baseline methods, confirming that Multi-Path Reasoning alone offers a strong foundation for visualization code generation. While full feedback-driven optimization leads to additional performance improvements, this result highlights that the primary strength of VisPath lies in its capacity to explore and leverage diverse reasoning trajectories. The framework remains effective and adaptable, even with minimal refinements, further validating importance of its core design.
@@ 4.3.3 Distinct Contribution of Visual Feedback
Incorporating rendered visual feedback improves the Executable Rate by $2\%-4\%$ and consistently boosts the Plot Score across both LLMs, as shown in Table 3. On GPT-4o mini, Plot Score increases from 64.82 to 66.12 (+1.30) and Executable Rate from 58% to 60% (+2 points). On Gemini 2.0 Flash, Plot Score rises from 57.68 to 59.37 (+1.69), and Executable Rate from 59% to 63% (+4 points). Despite the numerical gains are modest, the results demonstrate the unique value of structured visual evaluation. Visual feedback enables more refined and user-aligned outputs by capturing subtle rendering issues that may not affect executability, demonstrating its importance in the final synthesis stage.
LLM Variant Plot Score Executable Rate (%)
GPT-4o mini (w) feedback 66.12 60
(w) binary feedback 64.82 58
Gemini 2.0 Flash (w) feedback 59.37 63
(w) binary feedback 57.68 59
Table 3: Ablation results isolating the impact of rendered visual feedback. Comparing the full VisPath model using structured plot-based feedback with a binary-only feedback variant. Visual feedback provides consistent gains in both Plot Score and Executable Rate across LLMs.
@@ 5 Discussion
Our proposed VisPath framework substantially advances visualization code generation by addressing the core weaknesses of existing methods: limited interpretive flexibility and insufficient refinement. By employing Multi-Path Reasoning, VisPath explores diverse interpretations of user intent, which leads to more accurate visualizations, especially for ambiguous queries. Experimental results confirm its superiority: VisPath outperforms all baselines on both MatPlotBench and the Qwen-Agent Code Interpreter benchmark, with up to 9.14% improvement in Plot Score and 10% in Executable Rate. Ablation studies further validate VisPath’s design. First, increasing the number of reasoning paths enhances both visual quality and code executability. Second, even without visual feedback, Multi-Path Reasoning alone proves highly effective. Third, using structured plot-based feedback, rather than binary execution signals, significantly improves output alignment with user intent, confirming the value of our feedback-driven optimization module.
The figures provided in Appendix clearly illustrate VisPath’s robustness. [^4] The detailed cases are provided in Appendix B. In Case 1, it correctly centers conditional means using adjusted strip widths, unlike other methods that misalign pointplots or ignore key parameters. In Case 2, VisPath generates a proper polar bar chart with radial coordinates and legible labels, while others either fail to render or produce standard rectangular plots, with text often overlapping or unreadable. Case 3 highlights VisPath’s ability to handle compositional visualization requests: it correctly interprets the instruction to "visualize in 3 different ways" by generating a multi-subplot layout that includes line, scatter, and bar plots. Other approaches, by contrast, either collapsed all time series into a single chaotic plot or failed to differentiate the visualization modalities at all. These examples demonstrate how VisPath’s Multi-Path Reasoning and visual feedback lead to more precise and semantically aligned visualizations than current alternatives.
@@ 6 Conclusion
In this work, we present VisPath, a framework that leverages Multi-Path Reasoning and feedback-driven optimization to enhance automated visualization code generation. Unlike prior methods, our approach seamlessly combines Multi-Path Reasoning with feedback-driven optimization. Accurately capturing diverse user intents and refining generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark. By prioritizing adaptability, VisPath is uniquely positioned to handle ambiguous user queries through a combination of diverse reasoning paths and visual feedback integration. Future work could explore VisPath’s adaptability in more dynamic, real-world scenarios, further broadening its scope and practical utility in complex data analysis contexts.
7 Limitations
@@
Despite its effectiveness, the current framework relies on a limited feedback mechanism focused on query-code and query-plot alignment. While informative, these signals may overlook finer-grained elements essential to interpretability. Thus, future work could improve feedback depth by assessing individual plot components, such as readability and visual coherence, enabling more precise and refined visualization code generation. Moreover, while achieving strong performance, VisPath requires several rounds of agent interaction, including multi-path reasoning, execution, and feedback integration, which may introduce inefficiencies in certain use cases. Future work could explore ways to selectively identify the most promising reasoning paths early in the process, reducing redundant computation while preserving the benefits of diverse interpretation.
Appendix A Appendix A. Prompts Used
Prompt for Multi-Path Reasoning [System Prompt] According to the user query, expand and solidify the query into detailed instruction on how to write python code to fulfill the user query’s requirements. Import the appropriate libraries. Pinpoint the correct library functions to call and set each parameter in every function call accordingly. [User Prompt] Think step by step. Generate three distinct extended queries based on the given query. Ensure that you first analyze the given data description and create queries that align with the data. If no data description is provided, follow the initial query as is. You must follow the Python list format for the output. Do not modify the detailed instructions from the original user query. initial query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
Prompt for Code Generation [System Prompt] You are an expert in data visualization code generation. Think step by step and write the generated code in the format “‘python…“‘, where "…" represents the generated code. The code must end with ‘plt.show()‘. [User Prompt] Think step by step. Based on the user’s query and the provided data description, generate Python code using ‘matplotlib.pyplot‘ and ‘seaborn‘ to create the requested plot. Ensure that the code is formatted as “‘…“‘, where "…" represents the generated code. User query: {query} Data description: {datadescription}
Prompt for Visual Feedback [System Prompt] Given a code, a user query, and an image of the current plot, please determine whether the plot accurately follows the user query. Provide detailed instructions on how to enhance the plot using Python code. [User Prompt] Carefully analyze the provided Python code, the user query, and the plot image (if available) to assess whether the generated plot meets the user query requirements. If the plot image is missing, check the error message that occurred in the code. Compare the plot with the user query, highlight discrepancies, and provide clear, actionable instructions to modify the code. Additionally, suggest improvements for better visualization, focusing on clarity, readability, and alignment with the user’s objectives. Code: {code} User query: {oriquery}
Prompt for Synthesis [System Prompt] You are an expert on data visualization code judgement and aggregation. [User Prompt] Think step by step. Given the provided user query, data description, multiple data visualization codes generated for the query, and feedback for each code’s generated image. Your task is to: 1. Carefully review the user query and the data description. 2. Examine each version of the data visualization code along with the feedback provided for each version. 3. Synthesize the feedbacks for each code, user query insights, data description to create a final version of the code. 4. Your goal is to produce a final version of code that more effectively fulfills the user query by integrating the best elements from all versions and applying necessary corrections. User Query: {oriquery} Data Description: {datadescription} Code for aggregation with corresponding feedback: {codeforaggregation}
Prompt for Evaluation: MatplotBench You are an excellent judge at evaluating visualization plots between a model generated plot and the ground truth. You will be giving scores on how well it matches the ground truth plot. The generated plot will be given to you as the first figure. If the first figure is blank, that means the code failed to generate a figure. Another plot will be given to you as the second figure, which is the desired outcome of the user query, meaning it is the ground truth for you to reference. Please compare the two figures head to head and rate them. Suppose the second figure has a score of 100, rate the first figure on a scale from 0 to 100. Scoring should be carried out in the following aspect: 1. Plot correctness: Compare closely between the generated plot and the ground truth, the more resemblance the generated plot has compared to the ground truth, the higher the score. The score should be proportionate to the resemblance between the two plots. In some rare occurrence, see if the data points are generated randomly according to the query, if so, the generated plot may not perfectly match the ground truth, but it is correct nonetheless. Only rate the first figure, the second figure is only for reference. If the first figure is blank, that means the code failed to generate a figure. Give a score of 0 on the Plot correctness. After scoring from the above aspect, please give a final score. The final score is preceded by the [FINAL SCORE] token. For example [FINAL SCORE]: 40.
Prompt for Evaluation: Qwen-Agent Code Interpreter benchmark Please judge whether the image is consistent with the [Question] below, if it is consistent then reply "right", if not then reply "wrong". Question: {query}
Appendix B Appendix B. Case Study
Figure 4: Overview of the end-to-end process of visualizing conditional means alignment.
Figure 5: Overview of the end-to-end process of generating a polar bar chart.
Figure 6: Overview of the end-to-end process of visualizing time series data in three different ways.

### v3

@@ 4.1.2 Models Used
Large Language Models (LLMs)
For the code inference stage, we experiment with GPT-4o mini [1] and Gemini 2.0 Flash [26] to generate candidate visualization code from the reasoning paths. Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with previous work [37]. To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o [1] and Gemini 2.0 Flash [26] as our visualization feedback model, which provides high-quality reference assessments.
Vision-Language Models (VLMs)
In order to assess the visual quality and correctness of the rendered plots, we incorporate vision evaluation models into our framework. Specifically, GPT-4o [1] is employed for detailed plot evaluation in all evaluation tasks. This setup ensures the thorough evaluation of both the syntactic correctness of the code and the aesthetic quality of the resulting visualizations.
@@ 4.1.3 Evaluation Metrics
In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability. MatPlotBench [3] assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions. Qwen-Agent Code Interpreter benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/code_interpreter/README.md evaluates visualization models based on Visualization-Hard and Visualization-Easy. Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric. Previous studies showed that GPT-based VLM evaluations align well with human assessments [37], hence VLM was used for evaluation.
4.1.4 Baseline Methods
We compare VisPath against competitive baselines: (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses Chain-of-Thought (CoT) prompting to articulate its reasoning, while (3) Chat2VIS [19] employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent [37] first expands the query and then refines the code via a self-debugging loop with feedback. Moreover, our proposed framework VisPath generates three reasoning paths with corresponding visual feedback to refine the final output. For a fair comparison aligned with our experimental setting, MatPlotAgent is limited to three iterations, and uses a critique-based debugging loop.
Table 1: Performance comparison of various methods across different benchmarks. Visualization-Hard and Visualization-Easy refer to the Accuracy of Code Execution Results on different subsets of the Qwen-Agent Code Interpreter benchmark. Bold text indicates the best performance, underlined text indicates the second-best performance. denotes our proposed method.
@@ 4.1.4 Baseline Methods
CoT Prompting 61.95 50 57.50 40.00 48.75
Chat2VIS 56.98 53 59.36 36.50 47.93
MatPlotAgent 63.90 58 67.50 53.25 60.38
@@
CoT Prompting 53.56 61 40.00 63.89 51.95
Chat2VIS 54.89 55 59.36 56.50 57.93
MatPlotAgent 56.31 58 77.62 51.50 64.56
@@ 4.2 Experimental Analysis
VisPath is evaluated against four baselines, as shown in Table 1. Zero-shot prompting generates visualization code directly from natural language queries without intermediate reasoning. Despite its computational efficiency, this method frequently fails to handle ambiguity or under-specification, resulting in misaligned or incomplete outputs. On MatPlotBench (GPT-4o mini), it achieves a Plot Score of 62.38 and an Executable Rate of 53%. CoT prompting further introduces a single reasoning step to expose intermediate decisions and improve interpretability. However, on MatPlotBench, it slightly underperforms Zero-Shot in both Plot Score and Executable Rate, indicating reliance on a fixed reasoning path may reduce adaptability to diverse input structures.
Chat2VIS extends CoT prompting by adopting prefix templates to improve coherence and reduce ambiguity in user instructions. While effective for well-structured queries, its reliance on fixed templates limits adaptability to underspecified or unconventional requests. Such limitation is evident in its performance on MatPlotBench, where it achieves a Plot Score of 56.98 and an Executable Rate of 53%. Furthermore, MatPlotAgent incorporates query expansion and iterative self-debugging mechanisms to enhance robustness. While effective at correcting execution-level errors, its revisions are confined to localized adjustments and do not address higher-order semantic ambiguities.
@@
Evaluated across multiple benchmark settings, VisPath notably outperforms baselines, achieving up to 9.14 point gains in Plot Score and a 10% point increase in Executable Rate. These improvements demonstrate VisPath’s robustness in exploring diverse reasoning paths and refining outputs through visual feedback, which reduces semantic ambiguity and enhances execution reliability.
@@ 4.3 Ablation Study
To further examine the robustness and design choices of VisPath, we conduct a series of ablation experiments. Specifically, we analyze the following three aspects: (i) varying the number of generated reasoning paths, (ii) the effect of removing visual feedback during integration, and (iii) the contribution of visual feedback beyond binary executability.
@@ 4.3.1 Varying the Number of Reasoning Paths
To investigate the contribution of reasoning path diversity, we conducted ablation experiments by varying the number of generated reasoning paths $K$. In particular, we extended the range of $K$ from 2 to 8 to examine the effect of increased path on the overall performance of VisPath, as shown in Figure 2.
We observe a consistent pattern across all model and dataset combinations: performance improves as $K$ increases from 2 to 3, confirming that limited diversity $(K=2)$ often fails to capture nuanced interpretations of user queries. While $K=4$ achieves the highest executable rate on MatPlotBench with GPT-4o mini (62%), we further extend our analysis up to $K$ = 8 to comprehensively assess the impact of reasoning path diversity. However, beyond $K=4$, we observe diminishing returns and even performance degradation, which is likely due to noisy or redundant reasoning paths. While added diversity initially aids interpretation, excessive expansion burdens the integration process and reduces overall efficiency.
@@ 4.3.2 Robustness with a Simple Integration
Table 2: Performance comparison of VisPath with and without visual feedback. Results are reported on MatPlotBench (Plot Score, Executable Rate) and the average score on the Qwen-Agent Code Interpreter benchmark for two LLMs.
Model Setting MatPlotBench Qwen-Agent
Plot Score Exec. Rate (%) Avg.
GPT-4o mini w/o visual feedback 63.76 56 58.00
w/ visual feedback 66.12 60 63.96
Gemini 2.0 Flash w/o visual feedback 55.28 57 64.03
w/ visual feedback 59.37 63 68.98
Instead of refining each candidate visualization with VLM-based feedback, this approach aggregates three candidate codes: each derived from a distinct reasoning path, without intermediate corrections. This setup reduces computational overhead and execution time while preserving the benefits of interpretive diversity.
As shown in Table 2, even under this simplified configuration, VisPath outperforms all baseline methods, confirming that Multi-Path Reasoning alone offers a strong foundation for visualization code generation. While full feedback-driven optimization leads to additional performance improvements, this result highlights that the primary strength of VisPath lies in its capacity to explore and leverage diverse reasoning trajectories. The framework remains effective and adaptable, even with minimal refinements, further validating the importance of its core design.
Table 3: Ablation study isolating the impact of rendered visual feedback. Comparison between structured plot-based feedback and binary execution-only feedback in VisPath.
LLM Feedback Type Plot Score Exec. Rate (%)
GPT-4o mini Binary-only feedback 64.82 58
Visual feedback 66.12 60
Gemini 2.0 Flash Binary-only feedback 57.68 59
Visual feedback 59.37 63
@@ 4.3.3 Distinct Contribution of Visual Feedback
Incorporating rendered visual feedback improves the Executable Rate by $2\%-4\%$ and consistently boosts the Plot Score across both LLMs, as shown in Table 3. On GPT-4o mini, Plot Score increases from 64.82 to 66.12 (+1.30) and Executable Rate from 58% to 60% (+2 points). On Gemini 2.0 Flash, Plot Score rises from 57.68 to 59.37 (+1.69), and Executable Rate from 59% to 63% (+4 points). Despite the numerical gains being modest, the results demonstrate the unique value of structured visual evaluation. Visual feedback enables more refined and user-aligned outputs by capturing subtle rendering issues that may not affect executability, demonstrating its importance in the final synthesis stage.
Figure 3: (a) Scatter plot generation with explicit spatial and annotation constraints. User query: “Create a scatter plot of two distinct sets of random data, each containing 150 points. The first set (Group X) should be centered around (-2, -2) and visualized in blue, and the second set (Group Y) should be centered around (2, 2) and visualized in orange. Label each group at their respective centers with a round white box (…) ”
(b) Large-scale time-series visualization task. User query: “Visualize a large number of time series in three different ways.” This task evaluates each method’s ability to interpret an underspecified request, select appropriate plotting strategies, and compose multiple complementary visualizations for dense temporal data. GT denotes the ground-truth visualization.
@@ 5 Discussion
5.1 Effectiveness of Multi-Path Reasoning and Feedback-Driven Refinement
Our proposed VisPath framework substantially advances visualization code generation by addressing the core weaknesses of existing methods: limited interpretive flexibility and insufficient refinement. By employing Multi-Path Reasoning, VisPath explores diverse interpretations of user intent, which leads to more accurate visualizations, especially for ambiguous queries. Experimental results confirm its superiority: VisPath outperforms all baselines on both MatPlotBench and the Qwen-Agent Code Interpreter benchmark, with up to 9.14% improvement in Plot Score and 10% in Executable Rate.
Also, Fig. 3 compares VisPath with the baselines. VisPath matches the ground-truth structure more closely when GT is available, and still produces meaningful multi-view visualizations in Fig. 3(b) where no GT exists. The results further show that VisPath correctly follows layout constraints and handles vague time-series prompts, such as centering data groups and creating multiple views for large datasets. This suggests that the framework aligns more consistently with user intent in scenarios that remain challenging for other methods. Ablation studies further validate VisPath’s design. First, increasing the number of reasoning paths enhances both visual quality and code executability. Second, even without visual feedback, Multi-Path Reasoning alone proves highly effective. Third, using structured plot-based feedback, rather than binary execution signals, significantly improves output alignment with user intent, confirming the value of our feedback-driven optimization module.
5.2 Cost & Latency Trade-offs
To assess the latency impact of Multi-Path Reasoning and structured visual feedback, we compared the iteration counts of VisPath against the MatPlotAgent baseline on MatPlotBench and the Qwen benchmark. Iterations were measured across four categories: Query Expansion, Code Generation, Visual Feedback, and Editor. While MatPlotAgent uses dynamic iterations based on feedback, VisPath employs a fixed structure of $k=3$ reasoning paths for each query expansion.
Table 4: Iteration counts across components for MatPlotAgent and VisPath. Benchmarks differ in size: MatPlotBench has 100 rows, while the Qwen benchmark has 80 rows.
MatPlotBench (100 rows) Qwen (80 rows) MatPlotAgent VisPath MatPlotAgent VisPath Query Expansion 100 100 80 80 Code Generation 312 300 240 240 Visual Feedback 49 300 52 240 Editor 285 100 226 80 Total Iterations 746 800 598 640 Avg Iterations / Row 7.46 8.00 7.475 8.00
As shown in Table 4, MatPlotAgent averaged 7.46 iterations on MatPlotBench, while VisPath required 8. Importantly, VisPath incurs higher iteration counts for Visual Feedback but substantially fewer for the Editor stage (285 vs. 100 on MatPlotBench; 226 vs. 80 on the Qwen benchmark). Similarly, on the Qwen benchmark, MatPlotAgent averaged 7.475 iterations, with 8 for VisPath. This corresponds to a marginal increase of only 0.525–0.54 iterations, despite VisPath delivering substantial improvements in execution success and visual quality.
The ablation study (2) indicate that accuracy gains do not scale proportionally with $k$, suggesting that efficacy of VisPath is driven by its feedback-oriented design rather than brute-force iteration. By redistributing computational effort from editor retries to structured visual feedback, the system achieves meaningful performance without significant overhead. Since model calls directly impact latency and cost, $k=3$ provides the effective trade-off between robustness and efficiency for deployment beyond controlled benchmarks.
@@ 6 Conclusion
In this work, we present VisPath, a framework that leverages Multi-Path Reasoning and feedback-driven optimization to enhance automated visualization code generation. Unlike prior methods, our approach seamlessly combines Multi-Path Reasoning with feedback-driven optimization. By accurately capturing diverse user intents and iteratively refining the generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark. By prioritizing adaptability, VisPath is uniquely positioned to handle ambiguous user queries through a combination of diverse reasoning paths and visual feedback integration. Future work could explore VisPath’s adaptability in more dynamic, real-world scenarios, further broadening its scope and practical utility in complex data analysis contexts.
7 Limitation
@@
Despite its effectiveness, the current framework focuses on a feedback mechanism that assess query-code and query-plot alignment, which may overlook fine-grained elements essential to interpretability. Thus, future work could improve feedback depth by assessing individual plot components, such as readability and visual coherence, enabling more precise and refined visualization code generation.
Moreover, while achieving strong performance, VisPath requires several rounds of agent interaction, including multi-path reasoning, execution, and feedback integration, which may introduce inefficiencies in certain use cases. Future work could explore ways to selectively identify the most promising reasoning paths early in the process, reducing redundant computation while preserving the benefits of diverse interpretation.
Finally, our evaluation primarily focuses on quantitative metrics such as Plot Score and Executable Rate. While these metrics offer objective insights, they may not fully capture user-perceived quality, usability, or interpretability of the generated visualizations. Conducting user studies or expert reviews could provide complementary qualitative evidence and further validate the practical utility of VisPath.